# HI-SLAM2: Geometry-Aware Gaussian SLAM for Fast Monocular Scene Reconstruction

### v1

@@ IV-E Geometry Reconstruction Quality
TABLE III: Reconstruction evaluation on Replica dataset for implicit and explicit rgb methods. Ours surpass other methods especially large margin in accuracy.
@@
In Table III, we evaluate our geometry reconstruction results against recent NeRF-based and 3DGS-based methods on the Replica dataset, demonstrating superior performance in both accuracy and completeness metrics. As illustrated in Fig. 5, our method produces smoother reconstructions while preserving fine geometric details compared to GLORIE-SLAM [57] and HI-SLAM [36]. This is particularly evident in complex scene elements such as chair legs and shelf-mounted vases, where our results more closely match the ground truth. Qualitative comparisons on the ScanNet dataset (Fig. 6) further highlight our advantages over Splat-SLAM [33], showing more accurate geometry without floating artifacts and achieving better completeness. Additional qualitative results on the ScanNet++ dataset (Fig. 7) demonstrate our system’s capability to fully reconstruct challenging scenes, including low-texture surfaces like floors and walls. Notably, our reconstructions even capture glass windows that are missing in the laser scanner ground truth.
@@ IV-F Appearance Reconstruction Quality
TABLE IV: Rendering quality evaluations on Replica dataset for rgb and rgbd methods.
@@
TABLE V: Rendering quality evaluations on ScanNet dataset for rgb and rgbd methods.
@@
Tables IV and V present our rendering quality evaluation results. On the Replica dataset, our system significantly outperforms competing methods, achieving superior PSNR and LPIPS metrics. For the ScanNet dataset, we demonstrate better performance than RGB-D methods while matching the strong baseline of Splat-SLAM [33]. As detailed in our ablation study (Sec. IV-G), we could achieve even higher rendering quality by relaxing geometric constraints by normal loss, but this would compromise geometry accuracy. Instead, our system balances the trade-off between geometry and appearance quality.
@@ IV-G Ablation Study
TABLE VI: Depth accuracy of our final rendered depth compared to the prior depth aligned using different strategies, as well as BA depth with and without JDSA assistance, evaluated on the Replica dataset.
@@
TABLE VII: Ablation study on the progressive improvement in trajectory accuracy on the Replica dataset, averaged over 8 sequences. From left to right, each stage refines the pose estimation based on the previous stage.
@@
TABLE VIII: Ablation study on the impact of the different proposed modules on the reconstruction performance on the Replica dataset, averaged over 8 sequences.
@@
Monocular Prior Integration We first evaluate the depth estimation accuracy improved by our different modules on the Replica dataset, as depth accuracy is crucial for scene reconstruction quality. Table VI quantifies the effectiveness of different approaches. Comparing Prior (one scale) with Prior (scale grid) demonstrates that our grid-based scale alignment significantly outperforms the single-scale alignment from HI-SLAM [36], better addressing inherent scale distortions in monocular depth priors. Further analysis compares two depth estimation approaches: BA estimate (using BA alone) and BA with JDSA (using interleaved BA and JDSA optimization). The results confirm that incorporating JDSA with alternating optimization outperforms BA alone, validating our depth prior integration strategy. The final rendered depth from our Gaussian map achieves the highest accuracy, validating the effectiveness of our complete pipeline.
@@
Trajectory Accuracy Table VIII demonstrates the progressive improvement in pose estimation accuracy on the Replica dataset through our system pipeline. Starting with initial estimates from the online tracking module, accuracy is first enhanced through online PGBA based loop closing. A subsequent full BA further refines these results, with the final joint pose and 3DGS map refinement achieving the highest trajectory accuracy. This systematic improvement across stages validates the effectiveness of our hierarchical optimization approach.
Component Analysis To evaluate key design components, we conduct ablation studies by removing individual components. Table VIII confirms each module’s contribution to system performance. The grid-based scale alignment proves crucial, as its removal significantly degrades reconstruction accuracy. Similarly, the unbiased depth rendering enhances both rendering quality and geometric accuracy. While the normal loss slightly affects appearance metrics, it substantially improves geometry quality. The final joint pose and map refinement further enhances accuracy through improved global consistency.
@@
Our system achieves real-time performance with the online tracking, loop closing and mapping operating at 22 frames per second (FPS) on the Replica dataset, and takes only 12 seconds for offline refinement. On the ScanNet dataset, despite more rapid camera movements, it maintains robust performance online above 10 FPS. Fig. 9 illustrates the map size evolution for sequence Scene0000. While the Gaussian count initially grows during new area exploration, our efficient map pruning strategy stabilizes the map size by eliminating redundant Gaussians in revisited regions. This demonstrates our system’s scalability and efficiency in managing large-scale environments.
@@ V Conclusion
This article presents HI-SLAM2, a novel monocular SLAM system that achieves fast and accurate dense 3D scene reconstruction through four complementary modules. The online tracking module enhances depth and pose estimation by integrating depth priors with grid-based scale alignment, while parallel PGBA in the online loop closing module corrects pose and scale drift. Our mapping approach leverages 3D Gaussian splatting for compact scene representation, continuously refined during SLAM tracking. We enhance geometric consistency through monocular normal priors and unbiased ray-Gaussian intersection depth for splat-based rasterization. During offline refinement, we achieve high-fidelity reconstruction by incorporating exposure compensation and performing joint optimization of camera poses, 3DGS map, and exposure parameters. Extensive evaluations on challenging datasets demonstrate that HI-SLAM2 outperforms state-of-the-art methods in accuracy and completeness while maintaining superior runtime performance. Our system achieves high-quality geometry and appearance reconstruction without the typical trade-offs between the geometry and appearance in other methods. Future work will explore applications in challenging large-scale outdoor and dynamic environments.

### v2

@@ IV-E Geometry Reconstruction Quality
TABLE IV: Reconstruction evaluation on Replica dataset for implicit and explicit rgb methods. Ours surpass other methods especially large margin in accuracy.
@@
In Table IV, we evaluate our geometry reconstruction results against recent NeRF-based and 3DGS-based methods on the Replica dataset, demonstrating superior performance in both accuracy and completeness metrics. As illustrated in Fig. 5, our method produces smoother reconstructions while preserving fine geometric details compared to GLORIE-SLAM [59] and HI-SLAM [36]. This is particularly evident in complex scene elements such as chair legs and shelf-mounted vases, where our results more closely match the ground truth. Qualitative comparisons on the ScanNet dataset (Fig. 7) further highlight our advantages over Splat-SLAM [33], showing more accurate geometry without floating artifacts and achieving better completeness. Additional qualitative results on the ScanNet++ dataset (Fig. 9) demonstrate our system’s capability to fully reconstruct challenging scenes, including low-texture surfaces like floors and walls. Notably, our reconstructions even capture glass windows that are missing in the laser scanner ground truth.
@@ IV-F Appearance Reconstruction Quality
TABLE V: Rendering quality evaluations on Replica dataset for RGB and RGB-D methods.
@@
TABLE VI: Rendering quality evaluations on ScanNet dataset for rgb and rgbd methods.
@@
Tables III,  V and  VI present our rendering quality evaluation results. On the Replica dataset, our system significantly outperforms competing methods, achieving superior PSNR and LPIPS metrics. For the ScanNet dataset, we demonstrate better performance than RGB-D methods while matching the strong baseline of Splat-SLAM [33]. As detailed in our ablation study (Sec. IV-G), we could achieve even higher rendering quality by relaxing geometric constraints by normal loss, but this would compromise geometry accuracy. Instead, our system balances the trade-off between geometry and appearance quality.
Figure 8: Rendering quality comparison on the Waymo Open Dataset in unbounded outdoor scenes. Our method captures finer details of the driving environment, while other methods produce noticeably blurrier results.
On the Waymo Open dataset, Figure 8 shows qualitative comparisons of our rendered RGB images with those from other methods, and Table III presents the corresponding quantitative results. Our system achieves significantly higher rendering quality. This enhanced visual fidelity can be attributed to two key factors: firstly our superior tracking accuracy; and secondly the effective integration of depth and normal supervision in our mapping pipeline.
@@ IV-G Ablation Study
TABLE VII: Depth accuracy of our final rendered depth compared to the prior depth aligned using different strategies, as well as BA depth with and without JDSA assistance, evaluated on the Replica dataset.
@@
TABLE VIII: Ablation on the choice of depth prior models, evaluating rendering quality, reconstruction accuracy, and inference time on the Replica dataset.
PSNR  [dB] $\uparrow$ Acc  [cm] $\downarrow$ Comp  [cm] $\downarrow$ Comp Rat  [%] $\uparrow$ Runtime  [ms] $\downarrow$
Ours + Metric3D 38.59 1.60 3.58 85.11 61
Ours + DA V2 36.31 15.87 4.40 82.56 32
Ours + Omnidata 38.71 1.57 3.49 85.25 6
TABLE IX: Ablation study on the progressive improvement in trajectory accuracy on the Replica dataset, averaged over 8 sequences. From left to right, each stage refines the pose estimation based on the previous stage.
@@
TABLE X: Ablation study on the impact of the different proposed modules on the reconstruction performance on the Replica dataset, averaged over 8 sequences.
@@
Monocular Prior Integration We first evaluate the depth estimation accuracy improved by our different modules on the Replica dataset, as depth accuracy is crucial for scene reconstruction quality. Table VII quantifies the effectiveness of different approaches. Comparing Prior (one scale) with Prior (scale grid) demonstrates that our grid-based scale alignment significantly outperforms the single-scale alignment from HI-SLAM [36], better addressing inherent scale distortions in monocular depth priors. Further analysis compares two depth estimation approaches: BA estimate (using BA alone) and BA with JDSA (using interleaved BA and JDSA optimization). The results confirm that incorporating JDSA with alternating optimization outperforms BA alone, validating our depth prior integration strategy. The final rendered depth from our Gaussian map achieves the highest accuracy, validating the effectiveness of our complete pipeline. We further investigate the incorporation of alternative depth prior predictors, including Metric3D [77] and Depth Anything (DA) V2 [78], as shown in Table VIII. The results demonstrate that our method is compatible with these alternatives. However, our chosen prior, OmniData, provides a more efficient trade-off between performance and computational cost.
@@
Trajectory Accuracy Table IX demonstrates the progressive improvement in pose estimation accuracy on the Replica dataset through our system pipeline. Starting with initial estimates from the online tracking module, accuracy is first enhanced through online PGBA based loop closing. A subsequent full BA further refines these results, with the final joint pose and 3DGS map refinement achieving the highest trajectory accuracy. This systematic improvement across stages validates the effectiveness of our hierarchical optimization approach.
Component Analysis To evaluate key design components, we conduct ablation studies by removing individual components. Table X confirms each module’s contribution to system performance. The grid-based scale alignment proves crucial, as its removal significantly degrades reconstruction accuracy. Similarly, the unbiased depth rendering enhances both rendering quality and geometric accuracy, with qualitative comparison shown in Fig. 10. While the normal loss slightly affects appearance metrics, it substantially improves geometry quality. The final joint pose and map refinement further enhances accuracy through improved global consistency.
@@ IV-H Performance Analysis
Our system achieves real-time performance with the online tracking, loop closing and mapping operating at 22 frames per second (FPS) on the Replica dataset, and takes only 12 seconds for offline refinement. On the ScanNet dataset, despite more rapid camera movements, it maintains robust performance online above 10 FPS, with offline refinement taking a few minutes due to a higher number of optimization iterations. Fig. 11 illustrates the map size evolution for sequence Scene0000. While the Gaussian count initially grows during new area exploration, our efficient map pruning strategy stabilizes the map size by eliminating redundant Gaussians in revisited regions. This demonstrates our system’s scalability and efficiency in managing large-scale environments.
Figure 12: Evaluation on self-collected data with the robot navigating through a large factory hall. Our robot (a) is equipped with stereo cameras. Our system operates solely on the monocular input from the left camera and reconstructs the scene along with the estimated camera trajectory (b). The rendered color and depth maps (c) achieve comparable visual quality while exhibiting significantly better geometry and free of floaters that are present in the baseline DROID-SLAM + 3DGS.
IV-I Evaluation on Self-Collected Robot Data
We evaluate our system on a self-collected dataset captured by our robotic platform. Figure 12 shows the platform and the results in a large factory hall. The robot is equipped with stereo cameras. However, only the monocular input from the left camera is used for this experiment. The recorded sequence has intotal 4073 frames with a duration of 6 minutes and 52 seconds. We compare our rendering results with the baseline method DROID-SLAM + 3DGS, where the estimated camera poses and point cloud from DROID-SLAM serve as input and used to initialize the Gaussians. For our system, it completes online stage including tracking, loop closuing, and mapping in 5 minutes and 20 seconds, followed by 3 minutes and 32 seconds for the offline refinement stage. In contrast, the baseline requires 4 minutes and 24 seconds for DROID-SLAM and an additional 15 minutes and 2 seconds for 3DGS mapping. Overall, our method takes only about half the runtime of the baseline while achieving comparable rendering quality and significantly better geometry, without the floaters observed in the baseline results.
IV-J Extention to Semantic Reconstruction
Figure 13: Semantic reconstruction results: (a) outdoor scene from 100613 of Waymo Open dataset, (b) indoor scene from Room1 of the Replica dataset.
As an extension, we demonstrate the capability of our system for semantic scene reconstruction by incorporating 2D semantic information into the 3DGS representation. Each Gaussian primitive is augmented with semantic color channels in addition to its existing geometric and appearance attributes. These semantic channels can be efficiently rasterized to the image plane alongside color and depth, enabling simultaneous semantic colorization of the reconstructed scene. For semantic optimization, we maintain the same pipeline structure as our depth and pose optimization framework, with an additional L1 semantic RGB loss term that measures the absolute difference between rendered and ground truth semantic color maps. Following the evaluation protocol of [79], we assess our semantic reconstruction performance on the Replica dataset using mean Intersection over Union (mIoU) as the primary metric. We evaluate on four standard sequences to enable direct comparison with existing baseline methods.
TABLE XI: Semantic reconstruction results evaluated by mIoU metric on 4 sequences of the Replica dataset.
Method ro-0 ro-1 ro-2 of-0 Avg.mIoU[%$\uparrow$
RGB-D NIDS-SLAM[79] 82.45 84.08 76.99 85.94 82.37
DNS-SLAM[80] 88.32 84.90 81.20 84.66 84.77
SNI-SLAM[81] 88.42 87.43 86.16 87.63 87.41
SGS-SLAM[82] 92.95 92.91 92.10 92.90 92.72
RGB HI-SLAM[36] 74.93 79.55 80.90 71.53 76.72
Ours 90.27 92.80 91.11 92.45 91.65
Tab. XI presents quantitative results comparing our approach against both RGB-D and RGB-only semantic SLAM methods. Our system achieves competitive performance compared to state-of-the-art RGB-D methods, while significantly outperforming the RGB-only baseline HI-SLAM [36] by a margin of 14.93%. This substantial improvement can be attributed to two key factors: (1) our more accurate geometry reconstruction provides better surface boundaries for semantic label assignment, and (2) the explicit 3DGS representation allows for sharper semantic boundaries compared to implicit NeRF-based approaches that often struggle with object delineation in complex scenes or regions with small objects. Fig. 13 demonstrates qualitative results: indoor scene from the Replica dataset with ground truth semantic labels, and outdoor driving scene from the Waymo Open dataset using Mask2Former [83] predictions as semantic inputs.
@@ V Conclusion
This article presents HI-SLAM2, a novel monocular SLAM system that achieves fast and accurate dense 3D scene reconstruction through four complementary modules. The online tracking module enhances depth and pose estimation by integrating depth priors with grid-based scale alignment, while parallel PGBA in the online loop closing module corrects pose and scale drift. Our mapping approach leverages 3D Gaussian splatting for compact scene representation, continuously refined during SLAM tracking. We enhance geometric consistency through monocular normal priors and unbiased ray-Gaussian intersection depth for splat-based rasterization. During offline refinement, we achieve high-fidelity reconstruction by incorporating exposure compensation and performing joint optimization of camera poses, 3DGS map, and exposure parameters. Extensive evaluations on challenging datasets demonstrate that HI-SLAM2 outperforms state-of-the-art methods in accuracy and completeness while maintaining superior runtime performance. Our system achieves high-quality geometry and appearance reconstruction without the typical trade-offs observed in other methods.
Limitations: Our system has three main limitations: First, the current proximity-based loop closure detection shows limited robustness in the ETH3D dataset when encountering view occlusions and textureless regions, suggesting the need for learned feature-based place recognition. Second, in city-scale scenes, mapping quality can degrade due to limited optimization budget, indicating the need for submap optimization strategies. Third, the system assumes static environments. Incorporating dynamic object detection and tracking, along with motion segmentation, would enable robust operation in dynamic environments.